# Heterogeneity-Aware Federated Causal Inference Leveraging Effect-Measure Transportability

### v1

@@ 1 Introduction
Traditional approaches, such as meta-analysis, aggregate evidence from independent studies while accounting for study-level heterogeneity (Hartung et al.,, 2011). Extending beyond this classical framework, recent work in causal inference has sought to integrate data across heterogeneous sites while addressing violations of transportability, a notion often referred to under different but essentially equivalent terminologies, such as exchangeability, invariance and homogeneity. Examples include selective borrowing framework based on individual-level data (Gao et al.,, 2025), test-then-pool procedures (Yang et al.,, 2023; Gao and Yang,, 2023), and bias function modeling (Yang et al.,, 2025; Wu and Yang,, 2022; Cheng and Yang,, 2024). Some recent work has shifted toward federated settings, motivated by privacy constraints that prevent the pooling of individual-level data. Within this paradigm, robust estimators (Han et al.,, 2023; Xiong et al.,, 2023) and site-adaptive weighting schemes (Han et al.,, 2025; Liu et al.,, 2024) have been proposed to achieve efficient estimation by sharing only site-level aggregates.
@@
Recent studies have increasingly emphasized leveraging data heterogeneity across multiple sites to improve estimation of Average Treatment Effect (ATE) estimation, typically represented by the difference in means (Han et al.,, 2023; Xiong et al.,, 2023; Han et al.,, 2025). However, established clinical reporting standards, notably the CONSORT guidelines (Moher et al.,, 2010), emphasize the importance of presenting treatment effects through both absolute (e.g., risk difference) and relative (e.g., risk ratio) metrics to provide a more comprehensive assessment of treatment effects. The choice of scale has substantive implications: empirical evidence shows that effect framing significantly influences clinical decision-making (Naylor et al.,, 1992; Forrow et al.,, 1992), and treatment effect heterogeneity itself may be scale-dependent (Rothman et al.,, 2024). Motivated by these considerations, we study a broad class of causal measures, with a primary focus on the Risk Ratio (RR), to illustrate the our methodology and its properties.
@@
Some recent work addresses the integration of multiple data sources but often under strong transportability assumptions. Colnet et al., (2023) establishes identifiability of the causal estimand under a weaker condition by assuming the transportability of the treatment effect. In contrast to most approaches that assume transportability of the conditional outcome, we show that estimators based on this weaker assumption yield larger semiparametric bound but offer greater robustness due to fewer restrictions on the data likelihood function. Only a few studies have considered deriving efficient influence function (EIF) under this weaker assumption: Su et al., (2024) focuses solely on the mean difference, and Li, (2024) addresses the case with a single external data set. Our work fills these gaps by characterizing the EIF for a broader class of causal measures across multiple sites under this weaker assumption, an open problem explicitly noted in Su et al., (2024). Using the RR as an illustrative example, we derive its explicit EIF and highlight its desirable properties, including multiple robustness, rate robustness, and local efficiency. We also compare this EIF with the EIF derived under the stricter transportability assumption commonly used in the literature.
@@
As the complexity of data integration across multiple sites increases, addressing violations of key assumptions, such as transportability, has become crucial. Some recent work, for example Gao et al., (2025), considers individual selection method, but it is not developed in a federated setting. Most existing federated estimators employ a weighted framework to mitigate such violations (Liu et al.,, 2024; Han et al.,, 2025). However, weighting approaches often sacrifice statistical efficiency in finite samples, as the weights for incompatible sites cannot always be shrunk to zero. In contrast, directly applying the EIF-based estimator to the subset of sites that satisfy the necessary identification assumptions and conditions yields the smallest achievable semiparametric variance bound under this case. However, selecting the appropriate subset of sites becomes increasingly challenging as the number of sites grows, due to the exponential increase in possible combinations. To address this, we propose the Post-Federated Weighting Selection (PFWS) framework, which integrates federated weighting with a data-adaptive selection procedure to identify compatible sites. PFWS effectively alleviates the efficiency loss of weighted method under finite samples and attains the semiparaemtric efficiency bound asymptotically.
@@ 2 Notations, assumptions and preliminaries
To borrow the information from the source sites, additional assumptions are required. For causal measures satisfying Assumption 3, two identification strategies exist according to Colnet et al., (2023): (i) assuming transportability of the local effect measure, or (ii) assuming transportability of the conditional outcomes. These two strategies correspond to the following conditions:
@@ 3.1 Semiparametric efficient point estimators under ideal situation
where $\sigma^{2}_{a,k}(X)=\operatorname{var}(Y|A=a,S=k,X)$ for $k\in[[K]],\ a=0,1$, and $g^{(k)}(X)=\sigma^{2}_{1,k}(X)/\pi^{(k)}(X)+\sigma^{2}_{0,k}(X)\{\tau(X)\}^{2}/\{1-\pi^{(k)}(X)\}.$
@@
Similarly, as previously mentioned, we can obtain the estimator $\widehat{\psi}^{(0)}_{0}$ based on its EIF $\varphi_{0}^{(0)}$. The EIF for $\psi^{(0)}$, denoted $\varphi^{(0)}$, can be derived using the pathwise derivative. For instance, for RR, it can be calculated based on Lemma S2 in Jiang et al., (2022). The corresponding estimator for $\psi^{(0)}$ can be calculated as $\widehat{\psi}^{(0)}=m(\widehat{\psi}^{(0)}_{0},\widehat{\psi}^{(0)}_{1})$.
@@ 3.2 Theoretical properties
Assumptions 4 and 5 are standard regularity conditions used in M-estimation to achieve rate robustness (Van der Vaart,, 2000). Under Condition $1^{(1:K)}$ with Assumption 4, or under Condition $2^{(1:K)}$ with Assumption 5, the proposed estimators in $\S$3.1 can incorporate flexible methods for estimating the nuisance functions while still achieving parametric-rate consistency.
@@ 4 Heterogeneity-aware federated estimator
In practical setting, it is often unreasonable to assume that Assumptions 1(k) and 2(k), as well as Condition 1(k) hold across all source sites. In such cases, some source sites may provide relevant information for estimating the estimand, while other sites may not. In this section, we present an approach that combines information from both target site and source sites in a data-adaptive manner. We only assume that the identification assumptions hold for the target site, namely Assumptions $1^{(0)}$ and $2^{(0)}$. The algorithm is designed to detect source sites within which unconfoundedness, positivity assumptions, or Condition 1(k) may not hold, and to exclude these sites accordingly. This enables valid and efficient estimation, even in the presence of assumption violations and heterogeneity across sites. Our approach consists of two steps. In the first step, following Liu et al., (2024), we assign a weight for each site, which motivates the initial weighted estimator $\widehat{\psi}_{\text{fw}}^{(0)}$. In the second step, we propose a selection procedure based on the weights obtained in the first step. The result of the selection determines our final estimator, $\widehat{\psi}_{\text{fs}}^{(0)}$.
@@ 6 Real data analysis
Accurately estimating the effect of vasopressor administration on in-hospital mortality is critical in intensive care. We apply the proposed method to a real dataset from the eICU Collaborative Research Database (eICU-CRD) (Goldberger et al.,, 2000; Pollard et al.,, 2018, 2019), a multi-center ICU database containing de-identified health-related information from over 200,000 ICU admissions across the United States during 2014-2015.
@@ 7 Discussion
Rather than relying on selection to mitigate biases when borrowing external information, Yang et al., (2025) demonstrates that explicitly modeling and correcting the bias function can be an effective strategy, and extending this approach to federated settings is a promising direction for future work. Additionally, federated learning involving small sample sizes (Zhu et al.,, 2024) demands fundamentally different considerations, where methods such as randomization-based inference may be more appropriate.
@@
All technical details are provided in this Supplementary Material. Specifically, $\S$8 presents the general procedure for obtaining the EIF for various causal measures, along with the proofs of Theorems 1–7, $\S$9 gives the data generating process of our simulation studies, and $\S$10 presents additional diagnostics assessing the reasonableness of the sites selected by our method when applied to real data.
@@ 8.1 Efficient influence functions for various causal measures: a general procedure
To derive the EIF, we employ the parametric submodel approach of Bickel et al., (1993). Let $\{f_{t}(V):t\in\mathbb{R}\}$ denote a regular parametric submodel, with the true distribution corresponding to $t=0$, i.e., $f_{t}(V)|_{t=0}=f(V)$. Then, the tangent space $\mathcal{T}$ can be constructed by $\mathcal{T}=\mathcal{T}_{1}+\mathcal{T}_{2}+\mathcal{T}_{3}+\mathcal{T}_{4}$, where
@@
Back to the denominator of $\psi_{1}^{(0)}$, we can show that $\varphi_{1D}^{(0)}=I(S=0)-D_{1}^{(0)}$ is its efficient score. Then, the semiparametric efficient score for $\psi_{1}^{(0)}$ based on Lemma S2 in Jiang et al., (2022) is
@@ 8.2 Proof of Theorem 1
where $\sigma^{2}_{a,k}(X)=\operatorname{var}(Y|A=a,S=k,X)$ for $k\in[[K]],\ a=0,1$, and $g^{(k)}(X)=\sigma^{2}_{1,k}(X)/\pi^{(k)}(X)+\sigma^{2}_{0,k}(X)\{\tau(X)\}^{2}/\{1-\pi^{(k)}(X)\}.$
@@ 8.5 Proof of Theorem 4 and Theorem 6 (i)
    $\varphi_{1}^{(0)}$ belongs to a Donsker class (Van der Vaart,, 2000).
@@
where the third term (8) is $o_{p}(n^{-1/2})$ under 6 i). Even if the Donsker condition in Assumption 6 ii) is not met, the cross-fitting procedure in Chernozhukov et al., (2018) can be used to assure that (8) is negligible. We now show that the second term (7) is a small order term:
@@ 8.7 Proof of Theorem 7
The argument follows the oracle property proof structure for adaptive penalized estimators, inspired by the proof of Lemma 4 in Han et al., (2025). The key logic proceeds in two steps: (i) First, we establish that $\|\widetilde{\boldsymbol{w}}-\overline{\boldsymbol{w}}\|=O_{p}(n^{-1/2})$ by showing that the objective functions of the optimization problems defining $\widetilde{\boldsymbol{w}}$ and $\overline{\boldsymbol{w}}$ differ only by $O_{p}(n^{-1/2})$; (ii) Second, we show that $\widetilde{\boldsymbol{w}}=\widehat{\boldsymbol{w}}$ with high probability, which ensures the oracle selection set $\mathcal{S}$ can be correctly chosen asymptotically.

### v2

@@ 1 Introduction
Traditional approaches, such as meta-analysis, aggregate evidence from independent studies while accounting for study-level heterogeneity (hartung2011statistical). Extending beyond this classical framework, recent work in causal inference has sought to integrate data across heterogeneous sites while addressing violations of transportability, a notion often referred to under different but essentially equivalent terminologies, such as exchangeability, invariance and homogeneity. Examples include selective borrowing framework based on individual-level data (gao2025improving), test-then-pool procedures (yang2023elastic; gao2023pretest), and bias function modeling (yang2025data; wu2022integrative; cheng2024inference). Some recent work has shifted toward federated settings, motivated by privacy constraints that prevent the pooling of individual-level data. Within this paradigm, robust estimators (han2023multiply; xiong2023federated) and site-adaptive weighting schemes (han2025federated; liu2024multi) have been proposed to achieve efficient estimation by sharing only site-level aggregates.
@@
Recent studies have increasingly emphasized leveraging data heterogeneity across multiple sites to improve estimation of Average Treatment Effect (ATE) estimation, typically represented by the difference in means (han2023multiply; xiong2023federated; han2025federated). However, established clinical reporting standards, notably the CONSORT guidelines (moher2010consort), emphasize the importance of presenting treatment effects through both absolute (e.g., risk difference) and relative (e.g., risk ratio) metrics to provide a more comprehensive assessment of treatment effects. The choice of scale has substantive implications: empirical evidence shows that effect framing significantly influences clinical decision-making (naylor1992measured; forrow1992absolutely), and treatment effect heterogeneity itself may be scale-dependent (rothman2024epidemiology). Motivated by these considerations, we study a broad class of causal measures, with a primary focus on the Risk Ratio (RR), to illustrate the our methodology and its properties.
@@
Some recent work addresses the integration of multiple data sources but often under strong transportability assumptions. colnet2023risk establishes identifiability of the causal estimand under a weaker condition by assuming the transportability of the treatment effect. In contrast to most approaches that assume transportability of the conditional outcome, we show that estimators based on this weaker assumption yield larger semiparametric bound but offer greater robustness due to fewer restrictions on the data likelihood function. Only a few studies have considered deriving efficient influence function (EIF) under this weaker assumption: su2024efficient focuses solely on the mean difference, and li2024efficient addresses the case with a single external data set. Our work fills these gaps by characterizing the EIF for a broader class of causal measures across multiple sites under this weaker assumption, an open problem explicitly noted in su2024efficient. Using the RR as an illustrative example, we derive its explicit EIF and highlight its desirable properties, including multiple robustness, rate robustness, and local efficiency. We also compare this EIF with the EIF derived under the stricter transportability assumption commonly used in the literature.
@@
As the complexity of data integration across multiple sites increases, addressing violations of key assumptions, such as transportability, has become crucial. Some recent work, for example gao2025improving, considers individual selection method, but it is not developed in a federated setting. Most existing federated estimators employ a weighted framework to mitigate such violations (liu2024multi; han2025federated). However, weighting approaches often sacrifice statistical efficiency in finite samples, as the weights for incompatible sites cannot always be shrunk to zero. In contrast, directly applying the EIF-based estimator to the subset of sites that satisfy the necessary identification assumptions and conditions yields the smallest achievable semiparametric variance bound under this case. However, selecting the appropriate subset of sites becomes increasingly challenging as the number of sites grows, due to the exponential increase in possible combinations. To address this, we propose the Post-Federated Weighting Selection (PFWS) framework, which integrates federated weighting with a data-adaptive selection procedure to identify compatible sites. PFWS effectively alleviates the efficiency loss of weighted method under finite samples and attains the semiparaemtric efficiency bound asymptotically.
@@ 2 Notations, assumptions and preliminaries
To borrow the information from the source sites, additional assumptions are required. For causal measures satisfying Assumption 3, two identification strategies exist according to colnet2023risk: (i) assuming transportability of the local effect measure, or (ii) assuming transportability of the conditional outcomes. These two strategies correspond to the following conditions:
@@ 3.1 Semiparametric efficient point estimators under ideal situation
where $\sigma^{2}_{a,k}(X)=\operatorname{var}(Y|A=a,S=k,X)$ for $k\in[[K]],\ a=0,1$, and $g^{(k)}(X)=[\sigma^{2}_{1,k}(X)/\pi^{(k)}(X)+\sigma^{2}_{0,k}(X)\{\tau(X)\}^{2}/\{1-\pi^{(k)}(X)\}]/[\operatorname{pr}(S=k|X)\{\mu_{0}^{(k)}(X)\}^{2}].$
@@
Similarly, as previously mentioned, we can obtain the estimator $\widehat{\psi}^{(0)}_{0}$ based on its EIF $\varphi_{0}^{(0)}$. The EIF for $\psi^{(0)}$, denoted $\varphi^{(0)}$, can be derived using the pathwise derivative. For instance, for RR, it can be calculated based on Lemma S2 in jiang2022multiply. The corresponding estimator for $\psi^{(0)}$ can be calculated as $\widehat{\psi}^{(0)}=m(\widehat{\psi}^{(0)}_{0},\widehat{\psi}^{(0)}_{1})$.
@@ 3.2 Theoretical properties
Assumptions 4 and 5 are standard regularity conditions used in M-estimation to achieve rate robustness (van2000asymptotic). Under Condition $1^{(1:K)}$ with Assumption 4, or under Condition $2^{(1:K)}$ with Assumption 5, the proposed estimators in $\S$3.1 can incorporate flexible methods for estimating the nuisance functions while still achieving parametric-rate consistency.
@@ 4 Heterogeneity-aware federated estimator
In practical setting, it is often unreasonable to assume that Assumptions 1(k) and 2(k), as well as Condition 1(k) hold across all source sites. In such cases, some source sites may provide relevant information for estimating the estimand, while other sites may not. In this section, we present an approach that combines information from both target site and source sites in a data-adaptive manner. We only assume that the identification assumptions hold for the target site, namely Assumptions $1^{(0)}$ and $2^{(0)}$. The algorithm is designed to detect source sites within which unconfoundedness, positivity assumptions, or Condition 1(k) may not hold, and to exclude these sites accordingly. This enables valid and efficient estimation, even in the presence of assumption violations and heterogeneity across sites. Our approach consists of two steps. In the first step, following liu2024multi, we assign a weight for each site, which motivates the initial weighted estimator $\widehat{\psi}_{\text{fw}}^{(0)}$. In the second step, we propose a selection procedure based on the weights obtained in the first step. The result of the selection determines our final estimator, $\widehat{\psi}_{\text{fs}}^{(0)}$.
@@ 6 Real data analysis
Accurately estimating the effect of vasopressor administration on in-hospital mortality is critical in intensive care. We apply the proposed method to a real dataset from the eICU Collaborative Research Database (eICU-CRD) (goldberger2000physiobank; pollard2018eicu; pollard2019icu), a multi-center ICU database containing de-identified health-related information from over 200,000 ICU admissions across the United States during 2014-2015.
@@ 7 Discussion
Rather than relying on selection to mitigate biases when borrowing external information, yang2025data demonstrates that explicitly modeling and correcting the bias function can be an effective strategy, and extending this approach to federated settings is a promising direction for future work. Additionally, federated learning involving small sample sizes (zhu2024enhancing) demands fundamentally different considerations, where methods such as randomization-based inference may be more appropriate.
8 Acknowledgments
We thank Zehao Sun and Frank Eriksson for their careful review of an earlier version of this manuscript.
@@ 8 Acknowledgments
All technical details are provided in this Supplementary Material. Specifically, $\S$9 presents the general procedure for obtaining the EIF for various causal measures, along with the proofs of Theorems 1–7, $\S$10 gives the data generating process of our simulation studies, and $\S$11 presents additional diagnostics assessing the reasonableness of the sites selected by our method when applied to real data.
@@ 9.1 Efficient influence functions for various causal measures: a general procedure
To derive the EIF, we employ the parametric submodel approach of bickel1993efficient. Let $\{f_{t}(V):t\in\mathbb{R}\}$ denote a regular parametric submodel, with the true distribution corresponding to $t=0$, i.e., $f_{t}(V)|_{t=0}=f(V)$. Then, the tangent space $\mathcal{T}$ can be constructed by $\mathcal{T}=\mathcal{T}_{1}+\mathcal{T}_{2}+\mathcal{T}_{3}+\mathcal{T}_{4}$, where
@@
Back to the denominator of $\psi_{1}^{(0)}$, we can show that $\varphi_{1D}^{(0)}=I(S=0)-D_{1}^{(0)}$ is its efficient score. Then, the semiparametric efficient score for $\psi_{1}^{(0)}$ based on Lemma S2 in jiang2022multiply is
@@ 9.2 Proof of Theorem 1
where $\sigma^{2}_{a,k}(X)=\operatorname{var}(Y|A=a,S=k,X)$ for $k\in[[K]],\ a=0,1$, and $g^{(k)}(X)=[\sigma^{2}_{1,k}(X)/\pi^{(k)}(X)+\sigma^{2}_{0,k}(X)\{\tau(X)\}^{2}/\{1-\pi^{(k)}(X)\}]/[\operatorname{pr}(S=k|X)\{\mu_{0}^{(k)}(X)\}^{2}].$
@@ 9.5 Proof of Theorem 4 and Theorem 6 (i)
    $\varphi_{1}^{(0)}$ belongs to a Donsker class (van2000asymptotic).
@@
where the third term (8) is $o_{p}(n^{-1/2})$ under 6 i). Even if the Donsker condition in Assumption 6 ii) is not met, the cross-fitting procedure in chernozhukov2018double can be used to assure that (8) is negligible. We now show that the second term (7) is a small order term:
@@ 9.7 Proof of Theorem 7
The argument follows the oracle property proof structure for adaptive penalized estimators, inspired by the proof of Lemma 4 in han2025federated. The key logic proceeds in two steps: (i) First, we establish that $\|\widetilde{\boldsymbol{w}}-\overline{\boldsymbol{w}}\|=O_{p}(n^{-1/2})$ by showing that the objective functions of the optimization problems defining $\widetilde{\boldsymbol{w}}$ and $\overline{\boldsymbol{w}}$ differ only by $O_{p}(n^{-1/2})$; (ii) Second, we show that $\widetilde{\boldsymbol{w}}=\widehat{\boldsymbol{w}}$ with high probability, which ensures the oracle selection set $\mathcal{S}$ can be correctly chosen asymptotically.